# MAME: Multidimensional Adaptive Metamer Exploration with Human Perceptual Feedback

### v1

@@ 2.2.4 Selection of the Reference Stimuli Used in the Experiment
Lowest 20% conv1 1.2194 614
layer3 1.7333
avgpool 0.7942
@@
2.3 Generation of Target Images
This section describes the method for generating images at designated positions using gradient descent. The image modification process was performed by specifying the target layer (conv1, layer3, and avgpool), the ICA component of interest (components 1, 2, 3), the direction of change (positive or negative), and the magnitude of change (a positive scalar corresponding to the target value).
The image generation procedure is as follows:
  1. 1.
    Compute the ICA components of the reference image for the target layer as $\mathbf{y}_{\text{original}}=W\mathbf{x}_{\text{original}}$, where $\mathbf{x}_{\text{original}}$ is the vectorized gram matrix of the target layer.
  2. 2.
    Define the target component as $\mathrm{comp}$ and the target value as $t$. The modified ICA component $\mathbf{y}_{\text{target}}$ is given by:
    $\mathbf{y}_{\text{target}}=\mathbf{y}_{\text{original}}+t\mathbf{e}_{\text{comp}},$
    where $\mathbf{e}_{i}$ is a unit vector with 1 at the $i$-th component and 0 elsewhere. This operation modifies only the target component while keeping all others unchanged. If the change direction is negative, the unit vector is multiplied by $-1$.
  3. 3.
    In each iteration, compute the ICA components of the current image as $\mathbf{y}_{\text{current}}=W\mathbf{x}_{\text{current}}$.
  4. 4.
    Compute the mean squared error (MSE) between $\mathbf{y}_{\text{target}}$ and $\mathbf{y}_{\text{current}}$ as the loss function.
  5. 5.
Examples of image generation for the three layers and three components used in the experiment are shown in Fig.5. The eccentricity condition did not affect image generation; the same algorithm was used to generate images, which were then presented at different locations as a four-degree square image.
Figure 5: An example of the images used in the experiment, generated with three layers and three components each. The numbers (in red and blue) below the layer names represent the target ICA values used during image generation.
@@ 2.4 Feedback from Behavioral task
2.4.1 Participants and Experimental Environment
Eight participants took part in the experiment. All experiments were approved by the Ethics Committee of the University of Tokyo and conducted in accordance with the guidelines of the Declaration of Helsinki. Written informed consent was obtained from all participants.
@@ 2.4.1 Participants and Experimental Environment
2.4.2 ABX Task
The procedure of the ABX task is as follows (Fig. 6). In each trial, participants are presented with three stimuli in sequence: stimulus A, stimulus B, and stimulus X. The first two stimuli (A and B) are the reference and generated images in a random order. The third stimulus (X) is a test stimulus that is the same as A or B. Participants are required to identify whether X is the same as A or B.
The presentation position was 4°, 8°, or 12° from the center depending on the eccentricity condition. The size of the presentation image was 4° x 4° in the visual angle.
2.5 Exploration Using the Staircase Method
Based on the results of each ABX test, the target value was updated using the staircase method (2-up 1-down rule). Specifically, if a participant correctly answered two consecutive trials in the ABX task, the boundary of the human metameric space was narrowed by one step. Conversely, if the participant made an error in a single trial, the boundary was shifted in the direction of expansion. The threshold was estimated by averaging five reversals between correct and incorrect responses for each condition.
@@ 3 Results
The MAME framework was applied to each subject, and their metameric boundaries were obtained as ICA component values. Furthermore, analyses based on the RMS contrast of images and human contrast sensitivity were conducted, allowing the metameric boundaries to be computed using these indices.
@@ 3.1 Metameric Boundaries in ICA Components
The metameric boundaries represented by ICA components, obtained directly from the experiment, are shown in Table 2, Figure 7 and Figure 9. In Table 2 and Figure 7, the conditions of ICA components and direction were averaged within each participant; then, the mean thresholds were averaged across different participants. The horizontal axis of Figure 7 shows the layer conditions, and three eccentricities are plotted in each layer.
The results for eccentricity showed a gradual trend across conditions, with a particularly noticeable tendency for thresholds to increase as eccentricity increased in conv1 and layer3. To evaluate this trend in detail in terms of sensitivity to image contrast, the next section analyzes the results using RMS contrast
Figure 7: ICA threshold values at the metameric boundary. The mean and standard deviation across all subjects ($n=8$) are shown. The Conditional Min and Conditional Max in the figure represent the threshold search range, which depends on the step size and initial value for each layer. The upper bound is defined as the target threshold upper limit minus 0.4 times the step size, while the lower bound is the target threshold lower limit plus 0.4 times the step size.
Table 2: Mean threshold values computed by averaging across ICA components, search directions, and subjects for each layer and eccentricity. The standard deviation represents inter-subject variability ($n=8$).
@@
3.2 Metameric Boundaries Based on RMS Contrast
@@
$C_{\mathrm{RMS}}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(I_{i}-\langle I\rangle)^{2}}$
@@
The RMS contrast values exhibited a linear relationship with the ICA target values (Figures 8b, c, d). When visualizing the RMS contrast at the metameric boundary, it was observed that the values were particularly high in conv1 (Figure 8e). This suggests that, when measured using RMS contrast, human perception appears to be less sensitive to changes in conv1 compared to other layers.
Furthermore, the trend in eccentricity became more pronounced when analyzed using RMS contrast. Specifically, the slope of the threshold as a function of eccentricity was steepest in conv1, followed by layer3 and then avgpool, where it was more gradual.
Figure 8: (a) Examples of modulated images for each layer (displayed in grayscale). The target value $t$ determined based on the threshold for each layer. The rightmost column shows the difference images. (b, c, d) Relationship between ICA target values and the resulting RMS contrast for each layer (b: conv1, c: layer3, d: avgpool). For each target value, calculations were performed with $n=30$, randomly selecting ICA components and search directions. Error bars indicate standard deviation among difference images. (e) Mean and standard deviation of thresholds measured using RMS contrast. The average RMS contrast corresponding to the mean threshold values in Table 2 is shown, with standard deviation ($n=30$).
@@ 4.1 Summary of Main Findings
The present study proposed the MAME framework, an experimental framework to directly explore human metamers using online feedback by humans with weak supervision of artificial models. The effectiveness of this framework was demonstrated by successfully estimating 54-dimensional image metamer boundaries. Experimental results indicated that, among image generation layers, the threshold in conv1 was higher based on the magnitude of ICA thresholds and RMS contrast, suggesting that changes in this layer were less noticeable compared to others.
@@ 4.2 Alignments of Higher and Lower Visual Processing in Humans and Models
In our MAME framework, it is important to note that when an image is shifted even slightly in the model’s exploration direction from the reference image, the shifted image is no longer a model metamer for the model. This is because, in the MAME framework, exploration is conducted along specific directions in the model’s activation space. As a result, a single step in this space exceeds the model’s metameric boundary in the image dimension, as illustrated in Figure 1 (right). This notion suggests that in the present exploration, the lower the mean threshold for humans at a given hierarchical level, the closer the boundaries of human and model metamers, indicating a stronger alignment between human and model representations at that level. In this regard, the low threshold observed in avgpool suggests a closer functional alignment between the metameric spaces of humans and the model.
This finding is consistent with recent studies using robust CNNs (Feather et al.,, 2019, 2023; Gaziv et al.,, 2024). For instance, it has been shown that adversarial attacks on high-level activations can induce category modulations that are also perceptible to humans (Gaziv et al.,, 2024). Furthermore, the alignment of high-level features has been a topic of intense discussion in both visual and LLM literature. For example, the Platonic Representation Hypothesis (Huh et al.,, 2024) proposes that ’neural networks, trained with different objectives on different data and modalities, are converging to a shared statistical model of reality in their representation spaces.’ From this perspective, the strong alignment of high-level representations between humans and modern neural models, trained with various data, might be a straightforward consequence.
However, our experimental results also suggest a functional misalignment between models and humans in low-level processing. Given that convolutional neural networks (CNNs) are originally inspired by the information processing of simple and complex cells in the primary visual cortex (Fukushima,, 1980), this result may seem counterintuitive. Moreover, visualizing the first layer of a pre-trained CNN reveals the emergence of Gabor-like filters. However, a crucial factor is that feedforward models such as ResNet exhibit strong scale dependency in their information processing. The first-layer filters in ResNet50 have a high-resolution size of 7×7 pixels and lack low-frequency components. Additionally, while the correlation between filters is known to be important for visual processing (Freeman and Simoncelli,, 2011; Okazawa et al.,, 2015), not all correlations among early-layer features in CNNs (i.e., the Gram matrix) might be necessarily relevant. Some attempts to address such early-stage mismatches have been reported to contribute to the development of more generalizable models (Dapello et al.,, 2020). Therefore, based on our findings, reconsidering the early-stage architecture of models might be crucial for improving functional alignment between humans and models in the future.
@@ 4.3 Limitations and Broader Future Applications
The main limitation of our study is that the explored stimulus conditions in our experiment do not comprehensively cover all possible directions of human metamer space in terms of clarifying human functional processing. While We used Gram matrices to define metameric equivalence, many other possible conditions remain unexamined. For instance, metamers can be generated not only based on Gram matrices but also using activation differences. Conducting complementary experiments with such alternative conditions will be essential for further advancing our understanding of brain function.
Finally, it is noteworthy to discuss the generalizability of the proposed MAME framework. MAME is a versatile framework that extends beyond metamer exploration, as it allows for multidimensional modulation starting from a reference stimulus. The concept of providing feedback based on model-defined multidimensional features is analogous to Decoded Neurofeedback (Shibata et al.,, 2011), making it a promising candidate for various applications.
Moreover, since the framework can be executed in a web browser, it is free from hardware constraints. It can be applied even in environments with strong magnetic fields, such as fMRI, and is also suitable for crowdsourcing experiments. These advantages position MAME as a foundational tool for a wide range of applied research.

### v2

@@ 2.3 Generation of Target Images
Lowest 20% conv1 1.22 614
layer3 1.73
avgpool 0.79
@@ 2.4 Feedback from Behavioral task
As feedback in the present experiment, we employed an ABX task to measure perceptual similarity. While various forms of feedback could be considered—such as neural similarity or categorical judgments—our goal was to demonstrate the effectiveness of the novel MAME framework under specific conditions. In general, measuring perceptual appearance psychophysically is more time-consuming than categorical judgment in one trial. In contrast, computing the equivalence of neural or BOLD activations in neurophysiological or fMRI studies is similar time course to experiments for perceptual appearance, although the experimental cost is much higher in neuroscience experiments. The ABX task was chosen because it offers a comparable trial duration to neural similarity experiments, while being less costly to implement. This makes it a suitable choice for assessing the feasibility of the MAME framework in an online experimental environment.
@@
Eight participants took part in the experiment. The study was conducted with approval from the ethics committee of the research institution (detailed information will be disclosed after the double-blind peer-review process).
@@
The procedure of the ABX task is as follows (Fig. 5). In each trial, participants are presented with three stimuli in sequence: stimulus A, stimulus B, and stimulus X. The first two stimuli (A and B) are the reference and generated images in a random order. The third stimulus (X) is a test stimulus that is the same as A or B. Participants are required to identify whether X is the same as A or B. The presentation position was 4°, 8°, or 12° from the center depending on the eccentricity condition. The size of the presentation image was 4° x 4° in the visual angle.
Based on the results of each ABX test, the target value was updated using the staircase method (2-up 1-down rule). Specifically, if a participant correctly answered two consecutive trials in the ABX task, the boundary of the human metameric space was narrowed by one step. Conversely, if the participant made an error in a single trial, the boundary was shifted in the direction of expansion. Here, one step was defined as a fixed value for each layer, determined based on preliminary experiments: 10 for conv1, 0.3 for layer3, and 0.02 for avgpool. The threshold was estimated by averaging five reversals between correct and incorrect responses for each condition.
In summary, the experimental conditions consisted of three layers, three ICA components each explored in both positive and negative directions, and three levels of stimulus eccentricity (4°, 8°, and 12°). This resulted in a total of 54 conditions (3 × 3 × 2 × 3).
The experiment was conducted over two days for each participant. Each participant engaged in 15 experimental blocks, where each block consisted of 90 trials and took approximately 9–12 minutes to complete. In each block, the stimulus presentation position was fixed, and 18 conditions were presented five times each in a randomized order. Across the 15 blocks, three different stimulus presentation positions were used, with each position appearing in five blocks. The order of these positions was randomly determined at the start of the experiment. Following this protocol, each condition was tested 25 times using the ABX test.
@@ 3 Results
The MAME framework was applied to each subject, and their metameric boundaries were obtained as ICA component values. Furthermore, analyses based on the RMS contrast of images and Structural Similarity Index Measure (SSIM) were conducted, allowing the metameric boundaries to be computed using these indices.
@@ 3.1 Metameric Boundaries in ICA Components
The metameric boundaries represented by ICA components, obtained directly from the experiment, are shown in Table 2, Figure S5, and Figure S6. As the effects of the ICA component and direction conditions were small (Fig. S6), we first computed within-subject averages across these conditions. These mean values were averaged across different participants and shown in Table 2 and Figure S5.
Table 2: Mean threshold values computed by averaging across ICA components, search directions, and subjects for each layer and eccentricity. The standard deviation represents inter-subject variability, computed from the threshold values that were first averaged within each participant over components and directions. ($n=8$).
@@
The results for eccentricity showed a gradual trend across conditions, with a particularly noticeable tendency for thresholds to increase as eccentricity increased in conv1 and layer3. To evaluate this trend in detail in terms of sensitivity to image contrast, the next section analyzes the results using RMS contrast and SSIM.
3.2 Metameric Boundaries Based on RMS Contrast and SSIM
@@
$C_{\mathrm{R}}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(I_{i}-\langle\mathbf{I}\rangle)^{2}}$ (6)
@@
Figure 6: (a) Examples of threshold images for each layer (displayed in grayscale). The target value $t$ is determined based on the threshold for each layer. The rightmost column shows the difference images. (b) The RMS contrast of the difference image was computed. Images were generated with $n=30$ based on the average thresholds across participants from the behavioral experiment, and the mean and standard deviation of the RMS contrast of the difference image between the original and perturbed images are plotted. Lower values indicate smaller contrast differences. (c) Similarly, the SSIM between the original and perturbed images was computed, and the mean and standard deviation are plotted. Higher values indicate higher similarity between the original and perturbed images.
The RMS contrast values exhibited a linear relationship with the ICA target values. When visualizing the RMS contrast at the metameric boundary, it was observed that the values were particularly high in conv1 (Figure 6b). This suggests that, when measured using RMS contrast, human perception appears to be less sensitive to changes in conv1 compared to other layers.
Furthermore, the trend in eccentricity became more pronounced when analyzed using RMS contrast. Specifically, the slope of the threshold as a function of eccentricity was steepest in conv1, followed by layer3 and then avgpool, where it was more gradual. The finding that the effect of eccentricity diminishes at higher-level features has not been reported in previous studies and represents a novel contribution from the perspective of cognitive neuroscience.
Furthermore, these trends are not limited to comparisons based on RMS contrast, which reflects a global image statistic. We also evaluated perceptual sensitivity using the Structural Similarity Index (SSIM), which takes into account structural similarity between two images (Figure 6c). SSIM assesses local structural similarity within a sliding window. Consistent with the results based on RMS contrast, metameric images generated from the conv1 layer showed lower sensitivity, and the influence of eccentricity became progressively weaker as higher-level features were used.
@@ 4.1 Summary of Main Findings
The present study proposed the MAME framework, an experimental framework to directly explore human metamers using online feedback by humans with weak supervision of artificial models. The effectiveness of this framework was demonstrated by successfully estimating multi-dimensional image metamer boundaries. Experimental results indicated that, among image generation layers, the threshold in conv1 was higher based on the magnitude of ICA thresholds, RMS contrast, and SSIM, suggesting that changes in this layer were less noticeable compared to others. Furthermore, the effect of eccentricity on sensitivity was less pronounced for higher-level features.
To contextualize our findings, it is important to clarify what model metamers represent in the current study. In our MAME framework, when an image is shifted even slightly in the model’s exploration direction from the reference image, the shifted image is no longer a model metamer for the model. This is because exploration is carried out in the model’s activation space, where model metamers are defined as stimuli that are distinct in image space but equivalent in the target activation space. As a result, a single step in this space exceeds the model’s metameric boundary in image space.
This relationship implies that, in the current exploration paradigm, lower mean thresholds observed in human responses at a given hierarchical level indicate a closer alignment between the metamer boundaries of the human and the model. In this regard, the low threshold found in the avgpool layer suggests a closer functional alignment between the metameric spaces of humans and the model.
@@ 4.2 Alignments of Higher and Lower Visual Processing in Humans and Models
The finding related to the alignment of higher image features is consistent with recent studies using robust CNNs (Feather et al.,, 2019, 2023; Gaziv et al.,, 2024). For instance, it has been shown that adversarial attacks on high-level activations can induce category modulations that are also perceptible to humans (Gaziv et al.,, 2024). Furthermore, the alignment of high-level features has been a topic of intense discussion in both visual and LLM literature. For example, the Platonic Representation Hypothesis (Huh et al.,, 2024) proposes that ’neural networks, trained with different objectives on different data and modalities, are converging to a shared statistical model of reality in their representation spaces.’ From this perspective, the strong alignment of high-level representations between humans and modern neural models, trained with various data, might be a straightforward consequence.
However, our experimental results also suggest a functional misalignment between models and humans in low-level processing. Given that convolutional neural networks (CNNs) are originally inspired by the information processing of simple and complex cells in the primary visual cortex (Fukushima,, 1980), this result may seem counterintuitive. One reason is that the first-layer filters in ResNet50 have a high-resolution size of 7×7 pixels and lack low-frequency components. Additionally, while the correlation between filters is known to be important for visual processing (Freeman and Simoncelli,, 2011; Okazawa et al.,, 2015), not all correlations among early-layer features in CNNs (i.e., the Gram matrix) might be necessarily relevant. Some attempts to address such early-stage mismatches have been reported to contribute to the development of more generalizable models (Dapello et al.,, 2020). Considering our finding that higher-level features are less affected by eccentricity, reconsidering the early-stage architecture of models might be crucial for improving functional alignment between humans and models in the future.
@@ 4.3 Limitations and Broader Future Applications
The main limitation of our study is that the explored stimulus conditions in our experiment do not comprehensively cover all possible directions of human metamer space in terms of clarifying human functional processing. While we used Gram matrices to define metameric equivalence, activation differences could also be used. Conducting complementary experiments with such alternative conditions, involving a larger set of models and participants, will be essential for further advancing our understanding of brain function.
Finally, it is noteworthy to discuss the generalizability of the proposed MAME framework. Since the framework can be executed in a web browser, it is free from hardware constraints. It can be applied even in environments with strong magnetic fields, such as fMRI, and is also suitable for crowdsourcing experiments. These advantages position MAME as a foundational tool for a wide range of applied research.
5 Conclusions
In this study, we proposed the Multidimensional Adaptive Metamer Exploration (MAME) framework, which combines hierarchical neural network–based image generation with human-in-the-loop exploration, to directly map the multidimensional human metameric space. Our evaluation experiment demonstrated that MAME can measure this space within a single psychophysical experiment. The results further revealed that a recent human-aligned model exhibits a substantial functional mismatch with human processing for low-level features, despite closer alignment at higher representational stages. Because the user interface of MAME operates entirely within a web browser, it can be applied in fMRI environments and large-scale crowdsourcing experiments, making it a diverse tool for exploring metamers in human vision and for advancing interpretable AI.
6 Acknowledgments
We used ChatGPT (OpenAI, GPT-4) for English language editing. All content was reviewed and verified by the authors.